OpenStudy (anonymous):

Find the lateral area of a cylinder with base area of 49piin^2 and a height that is twice the radius of its base.

OpenStudy (wellingtonsm91):

First you need to find the height: Height = 2*radius But we know that: Area = pi*r^2 49piin^2 = pi*r^2 r = sqrt(49pin^2 / pi*r^2in) r = 7 in So height is: Height = 2 * 7 = 14 in Second- The lateral area (L): The lateral area (L) is: (circuference)*height (Circunference of the base) = 2*pi*r (inches) So: L = (2*pi*7)*14 (in^2) L = 87.96 in^2